# A covariant tapestry of linear GUP, metric-affine gravity, their Poincaré algebra and entropy bound

### 12 Jan 2024 v2

@@ Abstract
Several approaches to quantum gravity, including string theory, black hole physics, and quantum geometry, suggest the existence of a minimum measurable length. This concept, critical in theories like String Theory and Black Hole physics, poses significant questions about space-time’s fundamental structure. In String Theory, for instance, strings are theorized to not interact at distances smaller than their inherent size, suggesting a minimal measurable length. This assumption leads to the development of the Generalized Uncertainty Principle (GUP) [1, 2, 3, 4, 5, 6, 7, 8].
@@
The applications of GUP span a broad spectrum of fields and energy scales, ranging from atomic systems [14, 15], quantum optics [16, 17], and cold atoms [18, 19], to superconductivity and the quantum Hall effect [20], gravitational bar detectors [21], gravitational waves [22, 23], and even extending to gravitational decoherence [24], composite particles [25], astrophysical systems [26, 27], Baryon asymmetry [28], and condensed matter systems [29, 30]. Different theoretical approaches support linear GUP such as considering background coordinate time as a quantum degree of freedom coupled with mass-energy [31], discrete spacetime model of a graphene analog system [32]. GUP models are significant as they introduce modifications in state equations and microscopic variables[33, 34, 35, 36, 37, 38, 39]. These models predict potential quantum gravity effects observable at various scales, with theories often approximating a minimum length scale akin to the Planck length ($\ell_{Pl}=\sqrt{\frac{\hbar G}{c^{3}}}$), emphasizing quantum gravity’s role in the cosmic framework [40, 41, 42, 43, 44, 45, 46]. For a more comprehensive understanding of the GUP, its varied phenomenology, and experimental implications, there are detailed reviews available [47, 48, 49].
@@ III Covariant form of linear GUP and its symmetry
$[x_{\mu},p_{\nu}]=i\hbar\left[g_{\mu\nu}-\alpha\left(pg_{\mu\nu}+\frac{p_{\mu}p_{\nu}}{p}\right)\right]$ (14)
@@
$x_{\mu}$ $=$ $x_{0\mu}$ (15)
$p_{\mu}$ $=$ ${p_{0\mu}}(1-\alpha p_{0})$ (16)
where $x_{0\mu}$ and $p_{0\nu}$ satisfy the standard commutation relations: $[x_{0\mu},p_{0\nu}]=ig_{\mu\nu}\hbar$, and $p_{0}^{2}=p_{0\mu}p_{0}^{\mu}$. In this context, $p_{0\mu}$ represents the low-energy momentum expressed in position space as $p_{0\mu}=-i\frac{\partial}{\partial x_{0\mu}}$), whereas $p_{\mu}$ denotes the high-energy momentum, with $p_{0}$ being the norm of the $p_{\mu}$ vector. Using equations 15 and 16, the Lorentz algebra generator takes the following form:
$M^{\mu\nu}=\left(1-\alpha p_{0}\right)M_{0}^{\mu\nu}$ (17)
@@
$\left[x^{\mu},M^{\nu\rho}\right]=\left(1-\alpha p_{0}\right)~{}[x_{0}^{\mu},M_{0}^{\nu\rho}]-i\hbar\alpha\left(\frac{p_{0}^{\mu}}{p_{0}}+2\alpha p_{0}^{\mu}\right)M_{0}^{\nu\rho}$
@@
$\left[x^{\mu},M^{\nu\rho}\right]=-3i\hbar\alpha^{2}p_{0}^{\mu}M_{0}^{\nu\rho}$ (24)
@@
$\Gamma_{\mu\nu}^{\rho}\left[x^{\mu},M^{\nu}_{\rho}\right]\sim-i\alpha\hbar$ (25)
@@
These solutions are valid under the condition that $27~{}m^{2}c^{2}~{}\alpha^{2}\geq 1$. This bound may point to sub-Planckian physics [83, 84] if we set $m$ in the order of Planck mass $M_{Pl}$. In the absence of this condition, the solutions become complex, featuring imaginary components of the momentum. We notice here there exist three real solutions. One of these solutions presented by Eq. (27) is non-perturbative and remains unaffected by the particle’s mass. This particular solution is vital in the discretization of spacetime as evident in [9, 14, 12, 13]. The potential quantum gravity corrections indicated in the solutions presented in Eq. (28) and Eq. (29) might carry implications for phenomenological aspects in condensed matter, high energy, and astrophysical systems. Exploration of these implications is reserved for future research endeavors.
@@ IV Covariant GUP and Covariant Entropy bound
$MRc=\alpha^{\prime}\hbar$ (30)
@@
$\frac{S}{k_{B}}\leq\frac{A_{\text{3-Torus}}}{4\pi\ell_{Pl}^{2}}$ (37)
@@
$\frac{S}{k_{B}}\leq\frac{A_{\text{Sphere}}}{4\pi\ell_{Pl}^{2}}$ (38)
@@
$\Delta x\Delta p\geq\frac{S}{\pi k_{B}}~{}\frac{\hbar}{2}$ (39)
@@
$\Delta x\Delta p=\frac{A_{3-\text{Torus}}}{4\pi^{2}\ell_{Pl}^{2}}\frac{\hbar}{2}$ (40)
@@
$[x_{\mu},p_{\nu}]=i\hbar\left[g_{\mu\nu}-\alpha\left(pg_{\mu\nu}+\frac{p_{\mu}p_{\nu}}{p}\right)\right]=i\hbar~{}g_{\mu\nu}\frac{S}{\pi k_{B}}$ (41)
The Eqs. 40 and 41 provide a foundational basis for further exploration into the geometric interpretation of the linear GUP, which can be simply extended to various covariant minimal length theories [1, 2, 3, 4, 5, 6, 7, 8]. A figure showing the 3-Torus is given in Fig. (1).
@@
The torus is typically described using Riemannian geometry [105]. Its curved structure, which cannot be made flat without distortion, contrasts with the flat spacetime of Special Relativity where space is not curved by gravity or mass [106]. The depiction of the linear GUP through a 3-torus, grounded in Riemannian geometry, corroborates its equivalence with modified gravity[50]. The Eqs. 40 and 41 thus reflect a key relationship in Riemannian geometry, potentially leading to a broader understanding of quantum gravity.
@@ V Conclusions
In conclusion, we have developed a covariant version of a linear GUP model for Palatini-like theories of gravity. This is grounded in the recently proposed correspondence between metric-affine gravity and linear GUP at the phase space level. Consequently, we were able to derive the corresponding modified Poincaré algebra. This algebra exhibits distinct behavior, especially in the vicinity of nullity at the minimal length scale indicated by linear GUP, with the exception of the commutation relation between position and Lorentz generator. This non-vanishing commutation relation may point to a fixed background algebra in quantum gravity. A key element of our research is the adoption of 3-torus geometry for a covariant geometric visualization of GUP, which precisely describes geometrically the Bekenstein universal bound. Our approach retains the Schwarzschild radius ($r_{s}$) in its original form, diverging from the norm of substituting it with the radius ($R$) of the smallest sphere enclosing the system in the Bekenstein bound inequality. This leads to a re-envisioned covariant entropy bound, represented by the area of a torus with the inner radius determined by the Schwarzschild radius and the outer radius matching the smallest sphere enclosing the physical system. In our modest attempt, we propose a method that envisions the covariant entropy bound as the area of a torus, aiming to provide a more accurate geometric description when compared to Bousso’s covariant entropy bound. While Bousso’s bound, which is based on the area of the smallest sphere enclosing a physical system, offers valuable insights, our approach seeks to refine the geometric representation of the Bekenstein universal bound. We acknowledge that in certain cases, particularly with macroscopic systems, Bousso’s bound become larger than the Bekenstein universal bound, and our method is an effort to address this disparity. Additionally, in a tentative step forward, we suggest an equation that combines the linear covariant GUP with the 3-torus covariant entropy bound. This may open a new avenue for further exploration of quantum gravity. We hope to report on these in the future.
@@ Acknowledgments
In memory of Grzegorz Burda, you have become one of those stars we often spoke about during our teenage years.

### 05 Apr 2024 v3

@@ Abstract
Several approaches to quantum gravity, including string theory, black hole physics, and quantum geometry, suggest the existence of a minimum measurable length. This concept, critical in theories like String Theory and Black Hole physics, poses significant questions about space-time’s fundamental structure. In String Theory, for instance, strings are theorized to not interact at distances smaller than their inherent size, suggesting a minimal measurable length. This assumption leads to the development of the Generalized Uncertainty Principle (GUP) [1, 2, 3, 4, 5, 6, 7].
@@
The applications of GUP span a broad spectrum of fields and energy scales, ranging from atomic systems [13, 14], quantum optics [15, 16], and cold atoms [17, 18], to superconductivity and the quantum Hall effect [19], gravitational bar detectors [20], gravitational waves [21, 22, 23, 24], and even extending to gravitational decoherence [25], composite particles [26], astrophysical systems [27, 28], Baryon asymmetry [29], gravitational tests [30], and condensed matter systems [31, 32]. Different theoretical approaches support linear GUP such as considering background coordinate time as a quantum degree of freedom coupled with mass-energy [33], discrete spacetime model of a graphene analog system [34].
The study [35] demonstrates that linear GUP exhibits non-unitary properties, highlighting the significance of non-unitary processes in quantum mechanics. Non-unitary mechanics are essential for explaining various phenomena, including quantum decoherence, which describes the evolution of quantum states into classical states due to environmental interactions [36]; the dynamics of open quantum systems influenced by their surroundings [37]; the development of post-quantum theories integrating quantum mechanics with gravity [38]; the examination of PT-symmetric quantum mechanics, which studies non-Hermitian systems yielding real eigenvalues under parity-time symmetry [39]; and the exploration of quantum measurement and wave function collapse through models like GRW, which propose spontaneous localization to achieve measurement outcomes without observer intervention [40]. Furthermore, these non-unitary aspects contribute to the understanding of the information loss paradox in black holes, challenging the principle of unitary evolution and promoting research in quantum gravity [41].
In comparing linear and quadratic GUP models, [42] found quadratic GUP more consistent with full cosmological data, while linear GUP better explained late cosmological phenomena. Contrarily, [29] argued linear GUP could account for observed baryon asymmetry without necessitating a negative GUP parameter, unlike its quadratic counterpart that implies an unphysical imaginary minimal length. Essentially, while quadratic GUP necessitates a negative parameter in order to explain the Baryon asymmetry, leading to non-physical implications for the minimal measurable length, linear GUP supports a baryon asymmetry explanation with a physical, real minimal length. Baryon asymmetry, pivotal for cosmological models from Big Bang Nucleosynthesis to cosmic structure, underscores potential new physics beyond the Standard Model, including mechanisms like CP violation [43, 44, 45]. This could indicate the requirement to integrate both linear and quadratic versions of the GUP to thoroughly interpret full cosmological data and Baryon asymmetry, based on a real minimal measurable length.
GUP models are significant as they introduce modifications in state equations and microscopic variables [46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58]. These models predict potential quantum gravity effects observable at various scales, with theories often approximating a minimum length scale akin to the Planck length ($\ell_{Pl}=\sqrt{\frac{\hbar G}{c^{3}}}$), emphasizing quantum gravity’s role in the cosmic framework [59, 60, 61, 62, 63, 64, 65]. For a more comprehensive understanding of the GUP, its varied phenomenology, and experimental implications, there are detailed reviews available [66, 67, 68, 69, 70].
@@ III Covariant form of linear GUP and its symmetry
$[x_{\mu},p_{\nu}]=i\hbar\left[g_{\mu\nu}-\alpha\left(pg_{\mu\nu}+\frac{p_{\mu}p_{\nu}}{p}\right)\right],$ (14)
@@
$x_{\mu}$ $=$ $x_{0\mu},$ (15)
$p_{\mu}$ $=$ ${p_{0\mu}}(1-\alpha p_{0}),$ (16)
where $x_{0\mu}$ and $p_{0\nu}$ satisfy the standard commutation relations: $[x_{0\mu},p_{0\nu}]=ig_{\mu\nu}\hbar$, and $p_{0}^{2}=p_{0\mu}p_{0}^{\mu}$. In this context, $p_{0\mu}$ represents the low-energy momentum expressed in position space as $p_{0\mu}=-i\frac{\partial}{\partial x_{0\mu}}$), whereas $p_{\mu}$ denotes the high-energy momentum, with $p_{0}$ being the norm of the $p_{\mu}$ vector. In the work of Todorinov et al. [97], the Poincaré algebra was modified for the quadratic Generalized Uncertainty Principle (GUP), whereas in our research, we aim to develop the Poincaré algebra for the linear GUP. Using equations (15) and (16), the Lorentz algebra generator takes the following form:
$M^{\mu\nu}=\left(1-\alpha p_{0}\right)M_{0}^{\mu\nu}.$ (17)
@@
$\left[x^{\mu},M^{\nu\rho}\right]=\left(1-\alpha p_{0}\right)~{}[x_{0}^{\mu},M_{0}^{\nu\rho}]-i\hbar\alpha\left(\frac{p_{0}^{\mu}}{p_{0}}+2\alpha p_{0}^{\mu}\right)M_{0}^{\nu\rho}.$
@@
$\left[x^{\mu},M^{\nu\rho}\right]=-3i\hbar\alpha^{2}p_{0}^{\mu}M_{0}^{\nu\rho},$ (24)
@@
$\Gamma_{\mu\nu}^{\rho}\left[x^{\mu},M^{\nu}_{\rho}\right]\sim-i\alpha\hbar~{}.$ (25)
@@
These solutions are valid under the condition that $27~{}m^{2}c^{2}~{}\alpha^{2}\geq 1$. This bound may point to sub-Planckian physics [105, 106] if we set $m$ in the order of Planck mass $M_{Pl}$. In the absence of this condition, the solutions become complex, featuring imaginary components of the momentum. We notice here there exist three real solutions. One of these solutions presented by Eq. (27) is non-perturbative and remains unaffected by the particle’s mass. This particular solution is vital in the discretization of spacetime as evident in [8, 13, 11, 12]. The quantum gravity corrections highlighted in the solutions provided by Eq. (28) and Eq. (29) suggest possible impacts on a broader spectrum of physical systems. Specifically, these corrections could influence studies in condensed matter, high energy physics, and astrophysics, and extend to areas such as cosmology, quantum computing, and particle physics.
@@ IV Covariant GUP and Covariant Entropy bound
$MRc=\alpha^{\prime}\hbar~{}.$ (30)
@@
$\frac{S}{\pi k_{B}}\leq\frac{A_{\text{3-Torus}}}{4\pi^{2}\ell_{Pl}^{2}}~{},$ (37)
@@
$\frac{S}{\pi k_{B}}\leq\frac{A_{\text{Sphere}}}{4\pi^{2}\ell_{Pl}^{2}}~{},$ (38)
@@
$\Delta x\Delta p\geq\frac{S}{\pi k_{B}}~{}\frac{\hbar}{2}~{}.$ (39)
@@
$\Delta x\Delta p=\frac{A_{3-\text{Torus}}}{4\pi^{2}\ell_{Pl}^{2}}\frac{\hbar}{2}~{}.$ (40)
@@
$[x_{\mu},p_{\nu}]=i\hbar\left[g_{\mu\nu}-\alpha\left(pg_{\mu\nu}+\frac{p_{\mu}p_{\nu}}{p}\right)\right]=i\hbar~{}g_{\mu\nu}\frac{S}{\pi k_{B}}~{}.$ (41)
The Eqs. (40) and (41) provide a foundational basis for further exploration into the geometric interpretation of the linear GUP, which can be simply extended to various covariant minimal length theories [1, 2, 3, 46, 4, 5, 6, 7]. A figure showing the 3-Torus is given in Fig. (1).
@@
The torus is typically described using Riemannian geometry [127]. Its curved structure, which cannot be made flat without distortion, contrasts with the flat spacetime of Special Relativity where space is not curved by gravity or mass [128]. The linear GUP modeled through a 3-torus, embedded in Riemannian geometry, not only demonstrates compatibility with modified gravity theories but also illuminates a critical linkage in Riemannian geometry, as indicated by Eqs. (40) and (41). This geometric interpretation facilitates a broader comprehension of quantum gravity’s intricacies. The adoption of this model allows for an in-depth analysis of gravity’s role within quantum systems, contributing to a more nuanced understanding across disciplines such as quantum field theory, cosmology, material science, and quantum information. Thus, employing a 3-torus to represent GUP offers a geometric approach for exploring quantum gravity’s effects across diverse physical phenomena.
@@ V Conclusions
In conclusion, we have developed a covariant version of a linear GUP model for Palatini-like theories of gravity. This is grounded in the recently proposed correspondence between metric-affine gravity and linear GUP at the phase space level. Consequently, we were able to derive the corresponding modified Poincaré algebra. This algebra exhibits distinct behavior, especially in the vicinity of nullity at the minimal length scale indicated by linear GUP, except for the commutation relation between position and Lorentz generator. This non-vanishing commutation relation may point to a fixed background algebra in quantum gravity. A key element of our research is the adoption of 3-torus geometry for a covariant geometric visualization of GUP, which precisely describes geometrically the Bekenstein universal bound. Our approach retains the Schwarzschild radius ($r_{s}$) in its original form, diverging from the norm of substituting it with the radius ($R$) of the smallest sphere enclosing the system in the Bekenstein bound inequality. This leads to a re-envisioned covariant entropy bound, represented by the area of a torus with the inner radius determined by the Schwarzschild radius and the outer radius matching the smallest sphere enclosing the physical system. In our modest attempt, we propose a method that envisions the covariant entropy bound as the area of a torus, aiming to provide a more accurate geometric description when compared to Bousso’s covariant entropy bound. While Bousso’s bound, which is based on the area of the smallest sphere enclosing a physical system, offers valuable insights, our approach seeks to refine the geometric representation of the Bekenstein universal bound. We acknowledge that in certain cases, particularly with macroscopic systems, Bousso’s bound becomes larger than the Bekenstein universal bound, and our method is an effort to address this disparity. Additionally, in a tentative step forward, we suggest an equation that combines the linear covariant GUP with the 3-torus covariant entropy bound. This may open a new avenue for further exploration of quantum gravity. Moreover, the mentioned correspondence between modified gravity and linear GUP enables us to leverage tools developed by both communities for studying physical processes across large and small scales. For example, we can test the GUP using seismic data [61, 72], a method designed to constrain modified gravity theories. Conversely, the quantum-thermodynamic description of the non-relativistic limit of modified gravity allows us to test gravity models in tabletop experiments [129, 72]. Connecting (modified) gravity with mathematical structures describing the quantum properties of a given system, as provided in this paper, represents a fundamental step towards advancing us closer to formulating a quantum theory of gravity. Investigating these potential effects across such a diverse range of physical systems presents an exciting direction for future research. We hope to report on these in the future.
@@ Acknowledgments
In memory of Grzesiek Burda, you have become one of those stars we often spoke about during our teenage years.